OpenStudy (anonymous):

Does anyone know adding and subtracting rational numbers? -3 1/5 + (-6 1/2) a. -3 1/5 b. -3 7/10 c. -9 2/7 d. -9 7/10

OpenStudy (anonymous):

\[-3\frac{ 1 }{ 5 }=-\frac{ 3*5+1 }{ 5 }=-\frac{ 15+1 }{ 5 }=-\frac{ 16 }{ 5 }\] \[-6\frac{ 1 }{ 2 }=-\frac{ 6*2+1 }{ 2 }=-\frac{ 13 }{ 2 }\] \[-\frac{ 16 }{ 5 }+\frac{ -13 }{ 2 }=\frac{ -16*2+(-13*5) }{ 5*2 }=\frac{ -32-65 }{ 10 }=\frac{ -97 }{ 10}=-9\frac{ 7 }{ 10 }\]